OpenStudy (anonymous):

find the maximum range of domain,continuity and derivative of f(x)=arc sin(sin^2 x)+ arc tan sqrt(x)

OpenStudy (anonymous):

\[f(x)= Arc \sin (\sin ^{2}x) + Arc \tan \sqrt{x}\]

OpenStudy (anonymous):